May567:

only question #5 questions plz help

Vocaloid:

for 4 take the derivative of the function, set it equal to 0, and solve for x. that x value is the # of refridgerators at the maximum profit. plug that x-value back into the original function to find the profit 5a. if f(x) is your function and F(x) is the antiderivative...  applying this to 5a, take the integral, then evaluate F(4) - F(1) same logic with 5b

May567:

can you show me exactly the steps to do 5a? thanks @Vocaloid

Vocaloid:

integral of 6x is 3x^2 integral of 3/x is 3ln(x) (if you don't understand these, review your derivative and integration rules) so your antiderivative is F(x) = 3x^2 - 3ln(x) then evaluate F(4) - F(1) as these are the limits of integration

May567:

okay I got his part, what do you mean evaluate 1,4 do I pplug them in??

Vocaloid:

yes

May567:

3(4)^2 - 4 ln 4 = 53 3(1)^2 - 1 ln 1 = 3 53 - 3 = 50?? right? @Vocaloid

Vocaloid:

you don't change the 3 in front of the ln(x) 3x^2 - 3ln(x) ---> [3(4)^2 - 3ln(4)] - [3(1)^2 - 3ln(1)]

May567:

the answer is now equal to 49

May567:

but if you plug the whole thing in the calculator the answers need to be 40.8 @Vocaloid

Vocaloid:

you made a mistake in your arithmetic somewhere. check again.

May567:

so it should not be 49? im also using the calulator

Vocaloid:

if you're having trouble just use an online calculator wolframalpha.com and plug in this [3(4)^2 - 3ln(4)] - [3(1)^2 - 3ln(1)]

Vocaloid:

Sorry I gotta bail for a bit @imqwerty would you mind helping? we just need to wrap up 5a and start 5b, it's just a basic u-sub problem

May567:

thank you!

imqwerty:

@May567 did you remove the question? I can't find it